# EEG-based Graph-guided Domain Adaptation for Robust Cross-Session Emotion Recognition

### v1

@@ III-D Model Optimization
$\min_{\boldsymbol{S},\boldsymbol{A}}\sum_{j=1}^{n}\Big(\|\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\boldsymbol{A}^{T}\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big),$ (21)
@@
We denote $\boldsymbol{d}_{ij}={\|\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\boldsymbol{A}^{T}\boldsymbol{x}_{j}\|_{2}^{2}},$ and define $\boldsymbol{d}^{i}\in\mathbb{R}^{1\times n}$ as the vector whose $j$-th element is $d_{ij}$. problem (21) can be rewritten in the following row-wise form:
$\min_{\boldsymbol{S}}\frac{1}{2}\|\boldsymbol{s}^{i}+\frac{\boldsymbol{d}^{i}}{2\gamma}\|_{2}^{2},$ (22)
@@
Where, $\boldsymbol{s}^{i}$ denotes the $i$-th row of the similarity matrix $\boldsymbol{S}$, which is updated by solving (22). This update corresponds to a Euclidean projection onto the probability simplex and can be efficiently solved using the Lagrangian multiplier method combined with Newton’s method [20].
(2) $\boldsymbol{A}$ Optimziation: We can rewrite optimization problem (20) as:
$\min_{\boldsymbol{A}}\,\sum_{c=0}^{C}\operatorname{Tr}(\boldsymbol{A}^{T}(\alpha\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I})\boldsymbol{A})$ (23)
$\text{s.t. }\quad\boldsymbol{A}^{T}\boldsymbol{X}H\boldsymbol{X}^{T}\boldsymbol{A}=I,$
@@
$\mathcal{L}(\boldsymbol{A})=$ $\,\operatorname{Tr}\!\left(\boldsymbol{A}^{T}(\alpha\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\lambda\boldsymbol{I})\boldsymbol{A}\right)$ (24)
@@
where $\boldsymbol{\Phi}\in\mathbb{R}^{d\times d}$ is a diagonal matrix of Lagrange multipliers [21].
@@
$\left(\alpha\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{XL}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I}\right)\boldsymbol{A}$ (25)
$=\boldsymbol{XHX}^{T}\boldsymbol{A\Phi}.$
@@
$\boldsymbol{P}=\alpha\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{XL}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I},\qquad\boldsymbol{Q}=\boldsymbol{XHX}^{T},$
@@
The columns of $\boldsymbol{A}$ are the generalized eigenvectors, and the diagonal entries of $\boldsymbol{\Phi}$ are the corresponding eigenvalues. To minimize the objective, we select the $d_{f}$ eigenvectors associated with the smallest eigenvalues. Equivalently, $\boldsymbol{A}$ can be obtained from the $d_{f}$ eigenvectors of $\boldsymbol{Q}^{-1}\boldsymbol{P}$ corresponding to the smallest eigenvalues, assuming the null space of $\boldsymbol{X}$ has been removed so that $\boldsymbol{Q}$ is invertible.
After projecting the data into the learned subspace $\boldsymbol{A}$, a classifier (e.g., nearest neighbor) assign pseudo-labels to the target samples. These pseudo-labels are then used to update $\boldsymbol{M}_{c}$ and $\boldsymbol{S}$, refining class-wise distribution alignment. This procedure is repeated for a fixed number of iterations (typically 10–20). At each iteration, EGDAintegrates pseudo-label updating with an EM-like optimization strategy, progressively improving the target label estimates.
@@
Input: $\{\boldsymbol{X}_{l},\boldsymbol{Y}_{l}\}$, target $\boldsymbol{X}_{t}$; parameters $\alpha,\beta,\mu,\lambda,\gamma$; subspace dimension $d_{f}$.
Output: Predicted labels of the target samples $\boldsymbol{Y}_{u}$ and projection matrix $\boldsymbol{A}$;
@@
 Calculate the scatter matrix according to (7);Calculate the $\boldsymbol{M}_{0}$ according to (3) and the $\boldsymbol{M}_{c}$ according to (6).
@@
  Train a classifier $f$ on $(\boldsymbol{A}^{T}\boldsymbol{X}_{l},\boldsymbol{Y}_{l})$;
@@
until convergence
@@ IV Experiments
In this section, we evaluate the performance of the proposed EGDAmodel. We first describe the EEGdataset and the preprocessing steps. Then, we present the experimental setup and conduct two types of comparisons: (1) comparisons with existing methods evaluated on the same dataset, and (2) comparisons with representative methods that share similar modeling strategies. Finally, we introduce the performance measures used for evaluation. All experiments are conducted under a subject-dependent setting, where EEG data from the same subject are used for both training and testing across different sessions.
@@ IV-A Dataset Description
EEG signals were acquired using 62 electrodes arranged according to the international 10–20 system. The raw signals were preprocessed and segmented into 4-second EEG samples, which serve as the basic units for feature extraction. DEfeatures were then extracted under the assumption that EEG signals within each frequency band approximately follow a Gaussian distribution. DE features were computed from five standard frequency bands, namely Delta, Theta, Alpha, Beta, and Gamma. By concatenating DE features across all channels and frequency bands, each EEG sample is represented by a 310-dimensional feature vector. Owing to differences in video clip durations, each session contains approximately 830 EEG samples. .
@@ IV-B Overall Cross-Session Performance
As shown in Tables IIII, EGDAconsistently outperforms most existing methods whose performance has been previously reported on the SEED-IV dataset including GASDSL, OGSSL, SRAGL, GFIL, JSFA, CPTML and JAGP, achieving average accuracies of 81.22%, 80.15%, and 83.27% for the three transfer tasks, respectively. Although CPTML attains slightly higher accuracy in certain cases, it incurs substantially higher computational cost, which limits its applicability to large-scale or real-time scenarios. Overall, EGDAis ranked as the second-best method in terms of accuracy.
@@
Figure 1: Accuracies of different models for cross-session EEG emotion recognition
@@ IV-C Emotion-Specific Recognition
To further investigate per-class recognition, Fig. 2 provides a graphical mean confusion representation of 3 trials which reveals that the neutral state is recognized with the highest accuracy (87.83%), followed by happiness (76.52%). While, fear and sadness are more prone to misclassification, with frequent confusion observed between these two negative emotional states.
@@
subject GASDSL OGSSL SRAGL CPTML GFIL JSFA JAGPA EGDA
1 80.17 65.26 65.26 88.70 66.59 75.00 88.70 80.17
2 89.06 97.36 88.94 97.36 85.70 97.36 97.36 98.80
3 77.40 75.48 79.81 91.23 67.79 84.98 91.23 88.10
4 76.32 65.99 71.27 85.94 84.98 82.93 85.94 88.58
5 69.35 69.95 68.99 88.58 75.12 80.65 88.58 81.61
6 77.16 71.39 81.37 66.95 73.32 74.76 66.95 70.43
7 79.33 85.94 82.93 94.35 94.35 95.55 94.35 89.66
8 80.89 83.05 81.13 83.97 78.97 81.85 83.97 84.13
9 80.77 78.73 82.21 73.80 79.33 82.69 73.80 75.24
10 72.48 74.88 73.44 75.24 59.25 70.19 75.24 72.12
11 68.75 67.79 69.59 64.54 69.59 66.83 64.54 59.25
12 60.58 53.25 74.16 66.11 70.07 74.76 66.11 72.00
13 66.11 75.60 72.60 72.84 70.67 69.83 72.84 68.63
14 79.63 85.70 83.65 84.01 82.21 78.73 84.01 89.66
15 94.47 97.36 97.36 98.80 93.15 95.67 98.80 100
Average 77.08 76.51 78.18 82.16 75.33 80.79 79.78 81.22
@@
subject GASDSL OGSSL SRAGL CPTML GFIL JSFA JAGPA EGDA
1 75.84 80.54 73.48 84.67 74.09 69.83 79.44 84.67
2 87.23 91.85 93.43 88.69 80.54 64.84 68.61 57.18
3 83.45 66.42 67.03 74.21 75.06 78.47 71.05 65.09
4 79.83 79.68 80.05 93.07 86.74 88.69 80.78 83.21
5 73.84 52.43 65.21 85.40 66.18 85.40 76.52 80.66
6 76.64 85.89 81.39 84.79 82.12 94.28 88.81 91.36
7 88.81 91.85 96.35 80.54 79.20 96.84 86.74 89.17
8 79.56 80.90 85.04 94.89 81.02 77.13 92.82 92.94
9 75.43 77.98 79.56 80.54 70.19 75.55 75.91 80.54
10 73.48 73.97 75.30 73.51 68.61 77.37 69.34 76.28
11 78.59 78.83 80.66 85.89 71.53 83.94 72.51 78.83
12 65.94 55.84 74.94 73.51 56.57 75.67 58.34 66.30
13 74.09 64.48 73.11 63.02 57.91 63.75 64.72 62.41
14 83.73 84.31 89.66 89.54 85.28 91.73 89.42 97.69
15 85.52 91.24 93.07 92.85 90.27 92.58 97.32 95.86
Average 78.80 77.08 80.55 83.69 75.02 81.20 78.15 80.15
@@
NaturalHappyFearSad7.917.915.745.748.828.8287.8687.869.959.957.227.2210.8410.8470.1970.193.243.245.745.748.358.3576.5276.523.153.1514.3914.394.444.4475.9075.90
Figure 2: The average confusion graph of EGDAfor cross-session EEG emotion recognition.
@@ IV-D Critical Frequency Bands and Channels Identification
In this study, EEG feature vectors are constructed by concatenating DEfeatures extracted from multiple frequency bands and channels.
To evaluate the discriminative contribution of individual EEG features, we adopt the normalized $\ell_{2}$-norm of each row of the learned projection matrix $\boldsymbol{A}$ as a feature importance measure. Specifically, the importance of the $i$-th EEG feature is defined as:
@@
Since the feature representation is constructed by concatenating $62$ EEG channels across $5$ frequency bands, the importance vector $\boldsymbol{\theta}\in\mathbb{R}^{310}$ can be reshaped into a matrix $\boldsymbol{\Theta}\in\mathbb{R}^{62\times 5}$, where each column corresponds to a frequency band and each row corresponds to an EEG channel (Fig. 3). The overall importance of each frequency band is obtained by summing the elements within the corresponding column, while the importance of each EEG channel is computed by summing the elements within the corresponding row.
The results shown in Fig. 4 illustrate the average importance of each frequency band across three cross-session trials. Among all frequency bands, the Gamma band exhibits the highest contribution to emotion recognition, followed by the Alpha band.
At the channel level, Fig. 5 presents the ten most important EEG channels averaged across the three cross-session trials. These channels are primarily located in the prefrontal, central-parietal, and bilateral temporal regions, and demonstrate the highest discriminative values (Fig. 6). Furthermore, Fig. 7 highlights four EEG channels that are consistently identified as the most discriminative across all trials. These findings are consistent with previous studies, confirming the critical role of these brain regions in emotional processing [22].
@@
Figure 6: The topgraphical show of the average Importance channel of EGDAfor cross-session EEG emotion recognition.
@@
Figure 7: The four most important EEG channels selected for each cross-session experiment.

### v2

@@ III-C Model Optimization
$\min_{\boldsymbol{S}}\sum_{j=1}^{n}\Big(\|\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\boldsymbol{A}^{T}\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big),$ (21)
@@
We denote ${d}_{ij}={\|\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\boldsymbol{A}^{T}\boldsymbol{x}_{j}\|_{2}^{2}},$ and define $\boldsymbol{d}^{i}\in\mathbb{R}^{1\times n}$ as the vector whose $j$-th element is $d_{ij}$. Problem (21) can be rewritten in the following row-wise form:
$\min_{\boldsymbol{s}^{i}}\|\boldsymbol{s}^{i}+\frac{\boldsymbol{d}^{i}}{2\gamma}\|_{2}^{2},$ (22)
@@
where $\boldsymbol{s}^{i}$ denotes the $i$-th row of the similarity matrix $\boldsymbol{S}$, which is updated by solving (22). This update corresponds to a Euclidean projection onto the probability simplex and can be efficiently solved using the Lagrangian multiplier method combined with Newton’s method [20].
(2) $\boldsymbol{A}$ Optimization: We can rewrite optimization problem (20) as:
$\min_{\boldsymbol{A}}\ \operatorname{Tr}\!\left(\boldsymbol{A}^{T}\left(\begin{aligned} &\alpha\sum_{c=0}^{C}\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}\\ &+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I}\end{aligned}\right)\boldsymbol{A}\right),$ (23)
$\text{s.t. }\quad\boldsymbol{A}^{T}\boldsymbol{X}\boldsymbol{H}\boldsymbol{X}^{T}\boldsymbol{A}=\boldsymbol{I}.$
@@
$\mathcal{L}(\boldsymbol{A})=$ (24)
$\,\operatorname{Tr}\!\left(\boldsymbol{A}^{T}(\alpha\sum_{c=0}^{C}\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\lambda\boldsymbol{I})\boldsymbol{A}\right)$
@@
where $\boldsymbol{\Phi}\in\mathbb{R}^{d_{f}\times d_{f}}$ is a diagonal matrix of Lagrange multipliers [21].
@@
$\left(\alpha\sum_{c=0}^{C}\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I}\right)\boldsymbol{A}$ (25)
$=\boldsymbol{X}\boldsymbol{H}\boldsymbol{X}^{T}\boldsymbol{A}\boldsymbol{\Phi}.$
@@
$\boldsymbol{P}$ $=\alpha\sum_{c=0}^{C}\boldsymbol{X}\boldsymbol{M}_{c}\boldsymbol{X}^{T}+\beta\boldsymbol{S}_{w}+\mu\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}+\lambda\boldsymbol{I},$ (26)
$\boldsymbol{Q}$ $=\boldsymbol{X}\boldsymbol{H}\boldsymbol{X}^{T}.$
@@
The columns of $\boldsymbol{A}$ correspond to the generalized eigenvectors, and the diagonal entries of $\boldsymbol{\Phi}$ are the associated eigenvalues. To minimize the objective, we select the $d_{f}$ eigenvectors associated with the smallest eigenvalues. Equivalently, $\boldsymbol{A}$ can be obtained from the $d_{f}$ eigenvectors of $\boldsymbol{Q}^{-1}\boldsymbol{P}$ corresponding to the smallest eigenvalues, assuming that the null space of $\boldsymbol{X}$ has been removed, such that $\boldsymbol{Q}$ becomes invertible.
After projecting the data into the learned subspace $\boldsymbol{A}$, a 1-nearest neighbor (1-NN) classifier is employed to assign pseudo-labels to the target samples. These pseudo-labels are then used to update $\boldsymbol{M}_{c}$ and $\boldsymbol{S}$, refining class-wise distribution alignment. This procedure is iteratively performed for a fixed number of iterations, which is empirically determined. At each iteration, EGDA integrates pseudo-label updating with an EM-like optimization strategy, progressively improving the target label estimates.
@@
Input: $\{\boldsymbol{X}_{s},\boldsymbol{Y}_{s}\}$, target $\boldsymbol{X}_{t}$; parameters $\alpha,\beta,\mu,\lambda,\gamma$; subspace dimension $d_{f}$.
Output: Predicted labels of the target samples $\boldsymbol{Y}_{t}$ and projection matrix $\boldsymbol{A}$;
@@
 Calculate the scatter matrix according to (7);
 Calculate the $\boldsymbol{M}_{0}$ according to (3) and the $\boldsymbol{M}_{c}$ according to (6).
@@
  Train a classifier $f$ on $(\boldsymbol{A}^{T}\boldsymbol{X}_{s},\boldsymbol{Y}_{s})$;
@@
until the maximum number of iterations is reached
@@ IV Experiments
In this section, we evaluate the performance of the proposed EGDA model. We first describe the EEG dataset and the preprocessing steps. Then, we present the experimental setup and conduct two types of comparisons: (1) comparisons with existing methods evaluated on the same dataset, and (2) comparisons with representative methods that share similar modeling strategies. Finally, we introduce the performance measures used for evaluation. All experiments are conducted under a subject-dependent, cross-session setting, where EEG data from the same subject but different sessions are used as source and target domains, respectively.
@@ IV-A Dataset Description
EEG signals were acquired using 62 electrodes arranged according to the international 10–20 system. The raw signals were preprocessed and segmented into 4-second EEG samples, which serve as the basic units for feature extraction. differential entropy (DE) features were then extracted under the assumption that EEG signals within each frequency band approximately follow a Gaussian distribution. DE features were computed from five standard frequency bands, namely Delta, Theta, Alpha, Beta, and Gamma. By concatenating DE features across all channels and frequency bands, each EEG sample is represented by a 310-dimensional feature vector. Owing to differences in video clip durations, each session contains approximately 830 EEG samples.
@@ IV-B Overall Cross-Session Performance
As shown in Tables IIII, EGDA consistently outperforms most existing methods whose performance has been previously reported on the SEED-IV dataset including GASDSL, OGSSL, SRAGL, GFIL, JSFA, CPTML and JAGP, achieving average accuracies of 81.22%, 80.15%, and 83.27% for the three transfer tasks, respectively. Although CPTML attains slightly higher accuracy, it involves higher computational complexity, which may limit its applicability in large-scale or real-time scenarios. Overall, EGDA achieves the second-highest average accuracy among the compared methods.
@@
Figure 1: Comparison of recognition accuracies for different models in cross-session EEG emotion recognition
@@ IV-C Emotion-Specific Recognition
To further investigate class-wise recognition performance, Fig. 2 presents the averaged confusion matrix over three transfer tasks. The results show that the neutral emotional state is recognized with the highest accuracy (87.83%), followed by happiness (76.52%), whereas fear and sadness are more prone to misclassification.
@@
subject GASDSL OGSSL SRAGL CPTML GFIL JSFA JAGP EGDA
1 80.17 65.14 65.26 88.70 66.59 75.00 76.56 80.17
2 89.06 97.36 88.94 97.36 85.70 97.36 95.55 98.80
3 77.40 75.48 79.81 91.23 67.79 84.98 74.64 88.10
4 76.32 65.99 71.27 85.94 84.98 82.93 68.39 88.58
5 69.35 69.95 68.99 88.58 75.12 80.65 86.06 81.61
6 77.16 71.39 81.37 66.95 73.32 74.76 84.13 70.43
7 79.33 85.94 82.93 94.35 94.35 95.55 93.15 89.66
8 80.89 83.05 81.13 83.97 78.97 81.85 89.30 84.13
9 80.77 78.73 82.21 73.80 79.33 82.69 82.09 75.24
10 72.48 74.88 73.44 75.24 59.25 70.19 64.54 72.12
11 68.75 67.79 69.59 64.54 69.59 66.83 64.30 59.25
12 60.58 53.25 74.16 66.11 70.07 74.76 69.83 72.00
13 69.79 75.60 72.60 72.84 70.67 69.83 68.39 68.63
14 79.63 85.70 83.65 84.01 82.21 78.73 81.73 89.66
15 94.47 97.36 97.36 98.80 93.15 95.67 100 100
Average 77.08 76.51 78.18 82.16 75.33 80.79 79.91 81.22
@@
subject GASDSL OGSSL SRAGL CPTML GFIL JSFA JAGP EGDA
1 75.84 80.54 73.48 84.67 74.09 87.47 79.44 84.67
2 87.23 91.85 93.43 88.69 80.54 80.90 68.61 57.18
3 83.45 66.42 67.03 74.21 75.06 80.41 71.05 65.09
4 79.83 79.68 80.05 93.07 86.74 82.60 80.78 83.21
5 73.84 52.43 65.21 85.40 66.18 80.17 76.52 80.66
6 76.64 85.89 81.39 84.79 82.12 94.89 88.81 91.36
7 88.81 91.85 96.35 90.15 79.20 95.01 86.74 89.17
8 79.56 80.90 85.04 94.89 81.02 94.04 92.82 92.94
9 75.43 77.98 79.56 80.54 70.19 77.25 75.91 80.54
10 73.48 73.97 75.30 73.51 68.61 75.43 69.34 76.28
11 78.59 78.83 80.66 85.89 71.53 75.18 72.51 78.83
12 65.94 55.84 74.94 74.45 56.57 64.60 58.34 66.30
13 74.09 64.48 73.11 63.02 57.91 64.72 64.72 62.41
14 83.73 84.31 89.66 89.54 85.28 92.58 89.42 97.69
15 85.52 91.24 93.07 92.58 90.27 92.58 97.32 95.86
Average 78.80 77.08 80.55 83.69 75.02 82.52 78.15 80.15
@@
NeutralHappyFearSad7.917.915.745.748.828.8287.8687.869.959.957.227.2210.8410.8470.1970.193.243.245.745.748.358.3576.5276.523.153.1514.3914.394.444.4475.9075.90
Figure 2: The average confusion matrix of EGDA for cross-session EEG emotion recognition.
@@ IV-D Critical Frequency Bands and Channels Identification
In this study, EEG feature vectors are constructed by concatenating DE features extracted from multiple frequency bands and channels. To evaluate the discriminative contribution of individual EEG features, we adopt the normalized $\ell_{2}$-norm of each row of the learned projection matrix $\boldsymbol{A}$ as a feature importance measure. Specifically, the importance of the $i$-th EEG feature is defined as:
@@
Since the feature representation is constructed by concatenating $62$ EEG channels across $5$ frequency bands, the importance vector $\boldsymbol{\theta}\in\mathbb{R}^{310}$ can be reshaped into a matrix $\boldsymbol{\Theta}\in\mathbb{R}^{62\times 5}$, where each column corresponds to a frequency band and each row corresponds to an EEG channel (Fig. 3). The importance of each frequency band and EEG channel is obtained by summing the elements along the corresponding columns and rows, respectively.
The results shown in Fig. 4 illustrate the average importance of each frequency band across three cross-session tasks. Among all frequency bands, the Gamma band exhibits the highest contribution to emotion recognition, followed by the Alpha band.
At the channel level, Fig. 5 presents the ten most important EEG channels averaged across the three cross-session tasks. These channels are primarily distributed over the prefrontal, central-parietal, and bilateral temporal regions and exhibit the highest discriminative contributions, as illustrated in Fig. 6. Furthermore, Fig. 7 highlights four EEG channels that are consistently identified as the most discriminative across all trials. These findings are consistent with previous studies, confirming the critical role of these brain regions in emotional processing [22].
@@
Figure 6: Topographic map of the average channel importance obtained by EGDA for cross-session EEG emotion recognition.
@@
Figure 7: The four most important EEG channels selected for each cross-session task.